Ticket Number: HEA-NOC/20070228-2 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20070228 09:27 UTC Problem Start: 20070227 19:59 UTC
Ticket Update: 20070228 10:42 UTC Problem End:
NUIM's primary link went down at approx 19:59 last night. Flap seen on
other BT ATM circuits at the same time. NUIM not aware of any local
activity. HEAnet not aware of any planned activity in BT. NUIM backup link
is currently ISDN, so connectivity is degraded.
20070228 09:16 BS spoke to NUIM, opening ticket with BT.
20070228 09:27 BS fibre break last night at 19:59 - HEAnet not made aware
until contact made with NMC. BT sending across further details. BT ticket
20070228 10:38 BS BT NMC predicting fix by 4PM today. HEAnet investigating
use of new circuit with BT and NUIM, though it is likely that new circuit
is affected by the same break.
This ticket can be monitored at http://www.hea.net/tickets/20070228-2